# Synthetic data enables context-aware bioacoustic sound event detection

### v1

@@ 5 Experimental Evaluation
Table 2: F1 scores @0.3 IoU on FASD13. Methods marked with $\dagger$ were pre-trained using our generated data, rather than the data used in the original publication. Methods marked with involve no gradient updates at inference time.
5-shot, within-recording
Model AS CC GS HA HG HW JS KD MS PM RG RS RW Avg
BEATS+linear .350 .003 .056 .093 .242 .173 .028 .049 .462 .212 .732 .007 .316 .209
Protonet .356 .189 .156 .239 .038 .085 .136 .316 .590 .260 .000 .216 .393 .229
Protonet†∗ .305 .224 .151 .307 .023 .116 .166 .418 .536 .235 .121 .195 .342 .242
Transductive .299 .144 .002 .283 .020 .116 .279 .218 .569 .159 .089 .169 .048 .184
SCL .516 .333 .025 .438 .010 .255 .281 .263 .402 .237 .049 .219 .509 .272
SCL+finetuning .565 .341 .017 .467 .008 .382 .302 .381 .476 .327 .042 .285 .275 .298
SCL†∗ .545 .287 .024 .433 .008 .393 .243 .207 .429 .336 .038 .218 .228 .261
SCL+finetuning .571 .205 .030 .479 .005 .453 .132 .220 .516 .450 .050 .292 .223 .279
DRASDIC †∗ (ours) .645 .272 .593 .587 .144 .337 .099 .644 .783 .474 .092 .352 .764 .445
5-shot, cross-recording
Model AS CC GS HA HG HW JS KD MS PM RG RS RW Avg
BEATS+linear .114 .000 .139 .085 .063 .007 .002 .004 .349 .068 .605 .082 .133 .127
Protonet .404 .063 .084 .150 .025 .124 .132 .143 .570 .204 .000 .168 .281 .181
Protonet†∗ .255 .070 .100 .287 .012 .119 .084 .013 .413 .262 .000 .212 .228 .176
Transductive .042 .082 .001 .212 .015 .127 .352 .184 .536 .101 .108 .189 .043 .153
SCL .442 .157 .026 .254 .002 .218 .078 .155 .366 .216 .048 .263 .458 .207
SCL+finetuning .397 .187 .016 .344 .003 .170 .174 .178 .415 .340 .084 .331 .197 .218
SCL†∗ .469 .238 .032 .329 .005 .184 .169 .098 .437 .417 .044 .302 .193 .224
SCL+finetuning .408 .167 .037 .352 .005 .124 .193 .111 .461 .447 .096 .353 .209 .228
DRASDIC †∗ (ours) .500 .031 .641 .321 .091 .145 .004 .544 .650 .407 .058 .367 .689 .342
0-shot
Model AS CC GS HA HG HW JS KD MS PM RG RS RW Avg
BLED .505 .067 .138 .209 .060 .209 .739 .417 .518 .302 .338 .160 .177 .295
DRASD0S †∗ (ours) .163 .130 .259 .340 .116 .493 .000 .497 .549 .428 .082 .246 .645 .304
We evaluate models based on their ability to detect events after the fifth positive event in each recording of FASD13. Few-shot models (DRASDIC and its comparisons) have access to the audio and annotations up to the fifth positive event. Zero-shot methods use only high level features (the typical duration, high- and low-frequency, peak frequency, and SNR of target events). To set the value of the input features we measured these features on the first five positive events in a recording and took the median. Results for the validation set are in the Appendix. For all methods, we used validation set performance as the model selection criterion and only evaluated one final model version on FASD13.
@@ 5.1 Inference
For DRASDIC, for each of the five positive events in the support set, we take the $\operatorname{dur}_{s}$-seconds of labeled audio surrounding that event; this plus a $\operatorname{dur}_{q}$-second window of unlabeled audio forms one DRASDIC input. We average the detection logits for the unlabeled audio coming from all five of the inputs constructed this way. This is repeated for $\operatorname{dur}_{q}$-second windows across the entire query set. Inference with DRASD0S uses one forward pass through per query window. For both models, frames with predicted detection probability above a fixed threshold of 0.5 become positive detections. These are smoothed: detections separated by a gap of $\min(1,d/2)$ seconds are merged, and then detections lasting less than $\min(1/2,d/2)$ seconds are discarded. Here $d$ is the duration of the shortest event in the support set.
5.2 Metrics
Following [6], we evaluate models based on per-dataset detection F1. First, model detections are matched with annotated events. Detections and annotations with $\geq 0.3$ intersection-over-union (IoU) are identified as potential matches. Each detection is then assigned to at most one annotation, and vice-versa, using a graph matching algorithm [65]. Paired detections are true positives, unpaired detections are false positives, and unpaired annotations are false negatives.
@@ 5.3 Comparison methods
We compare DRASDIC with essentially all of the previously published methods we are aware of for FSBSED that contain publicly available implementations. The first, “BEATS+linear” is a simple supervised baseline which consists of a frozen BEATS encoder [66] and a final linear layer. Support audio is windowed (4 seconds, 50% overlap) and the final layer is trained for 100 epochs using cross-entropy loss. The initial learning rate of 0.01 (tuned using the validation set) is decayed to 0 using a cosine schedule. “Protonet” is the prototypical network from [8], which itself adapts [12]. “Transductive” [14] uses a CNN encoder that is updated using unlabeled audio from the query set. “SCL” applies the supervised contrastive learning method introduced by [13]. “SCL+finetuning”, also introduced by [13] extends this by using support audio to fine-tune the encoder that was pre-trained using the SCL method.
@@
For Protonet, SCL, and SCL+Finetuning, we train a version using the standard data from [6]. We also train a version using our generated scenes ($45e$ 4 scenes, each 40 seconds), which represents a 25×25$\times$25 × increase in data quantity over the data used for training the original models.
We compare DRASD0S with a band-limited energy detector (BLED), which is the only widely-used method we are aware of that performs zero-shot bioacoustic sound event detection based on high-level acoustic features. BLED detects events based on their within-band energy relative to a recording’s noise floor, and is included in Raven Pro, a standard software package used by bioacousticians [18]. Similar to DRASD0S, it requires the user to specify a target event’s frequency bounds, within-band SNR, and duration. For our experiments, we re-implemented BLED based on the Raven Pro documentation and tuned the detection threshold using the validation set. Detections produced by BEATS+Linear and BLED are smoothed in the same way as DRASDIC and DRASD0S.
@@
Figure 4: Qualitative results for few-shot task; events are in yellow. DRASDIC detects target sounds in dynamic environments (top two), but challenges include extremely low SNR (third), and the extended low-frequency drumming displays of ruffed grouse (bottom). Each spectrogram represents 10 seconds of audio.
5.4 Few-shot experiments
@@
We compare model performance on FASD13 (Table 2). For DRASDIC and DRASD0S, some datasets (JS, KD, MS, PB, PB24) contain events that are above the models’ 8kHz Nyquist frequency, or that are brief relative to the models’ 50Hz frame rate. For these, we give the model a 1/2-slowed version (1/6 for KD). For a fair comparison, we give other methods both the slowed and full-speed version of the data, and keep the version with the better score.
On the within-file version of FASD13, DRASDIC outperforms all the alternatives on 8 out of 13 datasets. Across datasets DRASDIC has an average F1 score of .147 over the next best model (49% improvement). For the cross-file version of FASD13, DRASDIC outperforms all others on 7 of 13 datasets and improves on the next best model by .114 F1 on average (50% improvement). The cross-file version is more difficult on average for all models; for DRASDIC the performance drops on average by .103 F1.
@@
Qualitatively, DRASDIC can detect a variety of target sounds, even in the presence of other sounds occupying the same frequency bands (Figure 4, top). Performance is strong across a variety of taxa and conditions. A failure case is for the JS dataset, which consists of jumping spider drumming. Here, the detection targets are specific drum types, and distinguishing between drum types relies partly on the rate of drumming. Our scene generator did not account for this type of information, but could be adapted to do so in the future. Other failure cases are in Figure 4, bottom.
@@
For the comparison methods we trained with our generated data, there was no clear performance increase. These methods, which adopt a CNN architecture, have fewer than $1/10$ the trainable parameters as DRASDIC, and likely do not have the capacity to take advantage of increased data scale.
5.5 Zero-shot experiments
Compared with BLED, DRASD0S performs better on 10 of 13 datasets in FASD13 (Table 2, bottom). Average performance is minorly better (.009 F1), but much of the contribution to the average F1 for BLED comes from its high performance on JS and AS. Compared with DRASDIC, DRASD0S performed worse with the exception of HW. It is possible that the additional information of high- and low-frequency bounds, which were not included in DRASDIC inputs, allowed DRASD0S to better locate low SNR vocalizations.
Table 3: Average F1 scores @0.3 IoU on FASD13 and validation datasets. All results are for our model (DRASDIC) trained with different data generation hyperparameters (top) or training/model ablations (bottom).
Data modification Avg (within) Avg (cross) Avg (val)
Reference .445 .342 .704
High homogeneity in events .429 .344 .630
Low homogeneity in events .444 .336 .606
High events / second .381 .292 .437
Low events / second .370 .298 .669
Only high SNR events .402 .300 .620
Only low SNR events .438 .333 .682
No pitch/time shifting .457 .376 .613
Ablation Avg (within) Avg (cross) Avg (val)
Reference .445 .342 .704
No curriculum learning .441 .362 .644
Shorter support (10 s) .406 .324 .618
No transformer .354 .261 .550
Reduced data (22 h) .130 .108 .088
Reduced data (220 h) .354 .253 .504
5.6 Additional experiments
@@
We investigated the impacts of adjusting the randomness governing our scene generation procedure (Table 3). We perturbed the level of homogeneity of target events in a scene, the typical rate of events, the loudness of events, and whether we apply pitch shifting augmentations (details in Appendix). Average performance is stable across some of these perturbations, but decreases when the randomness in event rate and event SNR is decreased. These parameters likely influence the level of diversity present across generated scenes more than the others.
@@
Eliminating random pitch shifts resulted in slightly better performance on FASD13. Designing a domain randomization strategy is an optimization problem [32], which we approached through a model selection criterion. This criterion did not produce the best model on the test set, aligning with the observation [8] that strong domain shifts between few-shot tasks present a challenge for FSBSED model development.
We performed ablations on our modeling and training approach (Table 3, bottom). Performance decreased when $\operatorname{dur}_{s}$ was reduced to 10 seconds. It also decreased when we removed the transformer and trained the remaining CNN audio encoder with a per-frame prototypical loss [67], a commonly adopted strategy for FSBSED [16]. Removing curriculum learning decreased validation performance but not performance on FASD13.
Finally, we investigated the effect of reducing the amount of training data available. We trained models with total duration of generated data equal to the duration of the train set adopted by other methods ($\approx 22$ hours), and also with 10×10$\times$10 × this amount. These performed poorly on FASD13.
@@ Conclusion
Appendix A FASD13 Summary
For all datasets, if there were overlapping bounding boxes corresponding to multiple events, we merged these bounding boxes into a single bounding box.
Anuraset:
Subset of twelve recordings from the strongly-labeled portion of AnuraSet [54]. The original study collected soundscape recordings from omni-directional microphones placed near four bodies of water in Brazil. The dataset was developed to improve automated classification of anuran vocalizations. Expert annotators identified the start- and end-times of vocalizations, for each of 42 different frog species. For our purpose, four recordings were selected for each of three frog species (Boana lundii, Leptodactylus latrans, Physalaemus albonotatus); annotations in these recordings that corresponded to other species were discarded.
Carrion Crow:
Set of ten hour-long recordings of carrion crows (Corvus corone) near León, Spain. Recordings were made using on-body recorders [68] attached to the tails of adult crows. Recordings were made through a study investigating communication and cooperative behavior in groups of crows. One expert annotator identified the start- and end-times of vocalizations of adult crows and socially-parasitic great spotted cuckoo (Clamator glandarius) chicks. For our purpose, five recordings were selected for each species; annotations in these recordings that corresponded to the other species were discarded. Vocalizations by crow chicks, such as begging calls, are considered as background sound. Crow vocalizations were marked as “Unknown” when it was difficult to discern the life stage of the vocalizing individual.
Gibbon:
Set of nine recordings from the Test split in [57]. Soundscape recordings were made in Hainan, China, using omni-directional microphones. The dataset was developed to improve monitoring of Hainan Gibbons (Nomascus hainanus), a critically endangered primate species. Expert annotators identified gibbon vocalizations, and annotated start- and end-times up to the closest second.
Gunshot:
Set of recordings from seven sites, taken from the Test split of the gunshot detection dataset presented in [55]. Recordings were made in forests in Gabon, using omni-directional microphones. The dataset was developed to investigate impacts of hunting on biodiversity in Gabon. Annotators marked the start- and end-times of gunshots. For our purpose, we collated all recordings from a single site into a single recording. Then, we discarded recordings which had fewer than seven detected gunshots.
Hawaiian Birds:
Subset of twelve recordings from the dataset presented in [56]. Recordings were made at a variety of locations in Hawaii, using omni-directional microphones. The recordings were collected for a variety of studies conducted by the Listening Observatory for Hawaiian Ecosystems at the University of Hawai‘i at Hilo. Expert annotators were asked to draw a spectrogram bounding box around each vocalization of 27 bird species present in Hawaii. Vocalizations separated by less than 0.5 seconds were allowed to be included in a single bounding box. For our purpose, four recordings were selected for each of three bird species (Chlorodrepanis virens, Myadestes obscurus, Pterodroma sandwichensis); annotations in these recordings that corresponded to other species were discarded.
Humpback
Subset of ten hour-long recordings from the dataset presented in [58]. Recordings were collected at sites in the North Pacific by bottom-mounted, omni-directional hydrophones. The dataset was developed to train a humpback whale (Megaptera novaeangliae) vocalization detector. We considered the “initial” audit portion of the data from this publication, in which experts annotated full recordings for humpback whale vocalizations and un-annotated time periods implicitly do not contain whale vocalizations. For our purpose, we selected ten recordings. The recordings were divided into 75-second chunks, and for each recording we discarded all subchunks which did not contain at least one whale vocalization.
Katydid
Subset of twelve recordings from the dataset presented in [59, 69]. Recordings were made in the forest canopy of Barro Colorado Island in Panamá, using an omni-directional microphone. The dataset was developed to quantify the calling activity of katydids (Tettigoniidae). Expert annotators identified the start- and end-times of katydid calls, for 24 species of katydids. For our purpose, we selected four recordings for each of three katydid species (Anaulacomera darwinii, Thamnobates subfalcata, Pristonotus tuberosus); annotations in these recordings that corresponded to other species were discarded. Additionally, following the original study, we only retained annotations where the annotators were able to identify a clear pulse structure.
Marmoset
Subset of ten recordings from the dataset presented in [70, 60]. Ten juvenile common marmosets (Callithrix jacchus) were each placed in a sound-proofed room, away from other individuals, and their spontaneous vocalizations were recorded using a cardioid microphone. The dataset was developed in order to investigate the use of deep learning for detecting and classifying marmoset vocalizations. Annotators identified the start- and end-time of each vocalization, and categorized each vocalization according to one of ten pre-defined call types. For our purpose, we selected one ten-minute recording from each individual. In five of these recordings, we retained annotations corresponding to the “Phee” call; in the other five recordings we retained annotations corresponding to the “Twitter” call. Similar call types (“Peep”, “Trillphee”, “Pheecry”, “TrillTwitter”, “PheeTwitter”) were re-labeled as “Unknown”, and the remaining annotated call types were discarded.
Powdermill
Recordings from the dataset presented in [61]. Four dawn chorus soundscapes were captured using omni-directional microphones at the Powdermill Nature Reserve, Pennsylvania, USA. The dataset was developed in order to provide a resource for research into automated bird sound classification and detection in soundscape recordings. Expert annotators marked the start- and end-times of vocalizations from 48 bird species. For our purpose, in two recordings we retained annotations of one bird species (Eastern Towhee, Pipilo erythrophthalmus), and in the other two recordings we retained annotations of a second bird species (Common Yellowthroat, Geothlypis trichas).
Rana Sierrae
Subset of recordings from the dataset presented in [63]. Underwater soundscapes were captured using omni-directional microphones placed in waterproof cases that were attached to the bottom of a lake in California’s Sierra Nevada mountains. The data were collected to characterize the vocal activity of a wild population of the endangered Sierra Nevada yellow-legged frog (Rana sierrae). For each vocalization, annotators marked its start- and stop-time, and classified it into one of five call types. For our purposes, we concatenated all recordings from each single day presented in the original dataset. This yielded seven recordings, corresponding to the seven days of recording. For four recordings, we retained annotations of one call type, “Primary vocalization”, and for the other three recordings, we retained annotations of the other call types “Frequency-modulated call”.
Right Whale
Subset of ten recordings from the dataset B presented in [71, 64]. Underwater soundscapes were recorded by hydrophones moored 5-50 meters above the bottom of the seafloor. The data were originally recorded as part of a study [64] documenting changes in the distribution of the endangered North Atlantic right whale (Eubalaena glacialis, NARW). Expert annotators manually midpoints of NARW upcalls. For our purpose, we extended each midpoint to a 1-second bounding box. The duration of this box was chosen based on the description of the NARW upcall in [71] as a “1-s, 100–200 Hz chirp with a 610 Hz bandwidth.”
Ruffed Grouse
Recordings from the dataset presented in [62]. Recordings were made using omni-directional microphones placed in regenerating timber harvests in Pennsylvania, USA. The dataset was developed to evaluate the performance of an automated method to detect ruffed grouse (Bonasa umbellus) drumming events. In the original study, five-minute clips were extracted from the original recordings, and annotators marked the start- and end-times of each drumming event. For our purpose, for each of the two months in the recording period (April and May, 2020), we concatenated all the recordings into a single audio file.
Spider
Four recordings provided by the Damian Elias lab. Male jumping spiders (Habronattus species) perform solid-borne acoustic displays in mating contexts. These displays were recorded using a laser vibrometer directed at the substrate on which spiders were standing. The dataset was collected as part of a study investigating signal evolution across the genus. Expert annotators labeled the start- and end-times of each spider signal, along with a signal-type category from a pre-defined list. For our purposes, for two files we retained labels signals labeled “thumping”; for the other two files we retained labels from signals labeled “knocking”.
Appendix B Data Generation Parameters
Sampling
For sampling target pseudovox, an event rate $r$ is drawn from $\{1,0.5,0.25,0.125,0.0625\}$ events per second. The number $n$ of pseudovox that will appear in a $\operatorname{dur}$-second scene is drawn from a Poisson distribution with rate parameter $r\times\operatorname{dur}$. To reduce the number of event-less scenes, we set $n=\max(n,1)$ with probability $1$ for support scenes and with probability $0.5$ for query scenes. A clustering level $k$ is drawn from $k\in K=\{\left\lfloor{M/128}\right\rfloor,\left\lfloor{M/64}\right\rfloor,\left\lfloor{M/32}\right\rfloor,\left\lfloor{M/16}\right\rfloor,\left\lfloor{M/8}\right\rfloor\},$ where $M$ is the total number of pseudovox. A cluster is drawn ar random from the $k$-means clustering of the pseudovox at this level, and $n$ pseudovox are drawn from this cluster. This process is repeated for distractor pseudovox.
For background audio, two background tracks are sampled, looped to the scene duration, and overlaid. When constructing support-query pairs, for the query scene these background tracks are different from the support scene background tracks with probability $p_{gen}=0.5$
Augmentation
With probability $0.2$, all target pseudovox are flipped in time. To apply pitch/time shifting, a resampling rate $\rho$ is drawn from $\{0.3,0.5,0.7,1,1,1,1.5,2\}$ and all target pseudovox are resampled from $16kHz$ to $\rho\times 16kHz.$ For amplitude augmentation, to simulate one or more individuals making sounds at different amplitudes, we construct a random Gaussian mixture model (GMM) with two components. Each component has a mean amplitude $\mu_{a}\sim\operatorname{Unif}(-12,7)$ dB and a standard deviation $\sigma_{a}\sim\operatorname{Unif}(0,5)$ dB. The weight of the second mixture component is drawn from $\{0,0,0,0,0,0.1,0.2,0.3,0.4,0.5\}$. The SNR of each target pseudovox is set based on draws from this GMM and set based on the RMS amplitude of the background audio and pseudovox. To add reverb, we convolve with a recorded room impulse response drawn from [72]. This process is repeated for distractor pseudovox. We apply only resampling augmentations to background audio.
Combination
To simulate one or more individuals making sounds at different rates, we construct a random GMM with two components that are used to sample time gaps between events in a scene. Each component has a mean $\mu_{t}\sim\operatorname{Unif}(0,30)$ seconds and a standard deviation $\sigma_{t}\sim\operatorname{Unif}(0,10)$ seconds. The weight of the second mixture component is drawn from $\{0,0,0,0,0,0.1,0.2,0.3,0.4,0.5\}$. Target pseudovox are added to the background audio, one-by-one, with timegaps between consecutive events sampled from this GMM. Events that extend past the duration of the scene are looped back to the beginning. This process is repeated for distractor pseudovox.
Variations for additional experiments
High homogeneity in events: Set cluster level $k=\lfloor{M/8}\rfloor$. Low homogeneity in events: Set cluster level $k=\lfloor{M/128}\rfloor$. High events / second: Set rate $r=1$ event per second Low events / second: Set rate $r=0.0625$ events per second Only high SNR events: SNR mean $\mu_{a}$ is drawn from $\operatorname{Unif}(2,7)$ dB. Only high SNR events: SNR mean $\mu_{a}$ is drawn from $\operatorname{Unif}(-12,-7)$ dB.
Appendix C Results on validation datasets
Table 4: F1 scores @0.3 IoU on validation datasets. Models marked with $\dagger$ were pre-trained using our generated data, rather than the data used in the original publication.
5-shot, within-recording
Model HB ME PB PB24 PW RD Avg
Frozen Encoder + Linear .839 .310 .067 .104 .668 .159 .358
Protonet .788 .597 .321 .492 .211 .359 .461
Protonet .775 .518 .480 .482 .165 .335 .459
Transductive .500 .173 .210 .342 .085 .143 .242
SCL .719 .691 .538 .688 .080 .368 .514
SCL+finetuning .779 .634 .577 .713 .077 .371 .525
SCL .578 .429 .374 .674 .090 .498 .440
SCL+finetuning .749 .486 .379 .659 .094 .353 .453
DRASDIC (ours) .659 .829 .657 .809 .738 .532 .704
0-shot
Model HB ME PB PB24 PW RD Avg
BLED .081 .469 .238 .588 .009 .458 .307
DRASD0S (ours) .000 .134 .424 .560 .070 .531 .287
Appendix D FASD13 visualizations
Figure 5: Example spectrograms from FASD13, part 1. Each row contains three spectrograms from one dataset. Positive events are highlighted.
Figure 6: Example spectrograms from FASD13, part 2. Each row contains three spectrograms from one dataset. Positive events are highlighted.
Figure 7: Example spectrograms from FASD13, part 3. Each row contains three spectrograms from one dataset. Positive events are highlighted.
Figure 8: Comparison of within- and cross-recording versions. For each support set, we computed the mean power mel-spectrum of A) the foreground (event) audio of the support set, B) the foreground audio of the within-recording query set, and C) the foreground audio of the cross-recording query set. The cosine similarity of A) and B) is on the horizontal axis, and of A) and C) on the vertical axis. 79% of support foreground audio is more similar to the within-recording query audio (i.e. below the line of slope=1) than to the cross-recording query audio. A similar trend holds for background audio: 69% of support background audio is more similar to the within-recording query audio than to the cross-recording query audio.

### v2

@@ 5 Experimental Evaluation
Table 2: F1 scores @0.3 IoU on FASD13. Methods marked with $\dagger$ were pre-trained using our generated data, rather than the data used in the original publication. Methods marked with involve no gradient updates at inference time. The second column gives the number of model parameters, and the final column gives the average F1 score across the six validation datasets from [6].
Model Params AS $C$ GS HA HG HW JS KD MS PM RG RS RW Avg Val
BEATS+linear 90M .350 .003 .056 .093 .242 .173 .028 .049 .462 .212 .732 .007 .316 .209 .358
AVES+linear 90M .586 .059 .500 .374 .207 .366 .026 .673 .831 .291 .529 .303 .494 .403 .565
Protonet 0.7M .356 .189 .156 .239 .038 .085 .136 .316 .590 .260 .000 .216 .393 .229 .461
Protonet†∗ 0.7M .305 .224 .151 .307 .023 .116 .166 .418 .536 .235 .121 .195 .342 .242 .459
Transductive 0.5M .299 .144 .002 .283 .020 .116 .279 .218 .569 .159 .089 .169 .048 .184 .242
SCL 7.2M .516 .333 .025 .438 .010 .255 .281 .263 .402 .237 .049 .219 .509 .272 .514
SCL+finetuning 7.2M .565 .341 .017 .467 .008 .382 .302 .381 .476 .327 .042 .285 .275 .298 .525
SCL†∗ 7.2M .545 .287 .024 .433 .008 .393 .243 .207 .429 .336 .038 .218 .228 .261 .440
SCL+finetuning 7.2M .571 .205 .030 .479 .005 .453 .132 .220 .516 .450 .050 .292 .223 .279 .453
DRASDIC †∗ (ours) 116M .645 .272 .593 .587 .144 .337 .099 .644 .783 .474 .092 .352 .764 .445 .704
We evaluate models based on their ability to detect events after the $N=5^{th}$ positive event in each recording of FASD13, using F1@0.3 IoU as described in [4]. We used performance on the validation set from [6] to select a final model to evaluate on FASD13.
@@ 5.1 Inference
For DRASDIC, we form predictions by windowing the audio in each recording, making multiple predictions for each window by prompting the model multiple times, and averaging these predictions. In detail, for a fixed $\operatorname{dur}_{q}$-second window of the query set, we prompt the model $N=5$ times and average the frame-wise predictions produced by these five prompts. The support set for the $i^{th}$ prompt ($i\in[1,5]$) is the $\operatorname{dur}_{s}$-seconds of support audio centered at the $i^{th}$ positive event in the support set (together with the binary detection mask). This procedure is repeated for $\operatorname{dur}_{q}$-second windows across the entire query set. Frames with predicted detection probability above a fixed threshold of 0.5 become positive detections. These are smoothed: detections separated by a gap of $\min(1,d/2)$ seconds are merged, and then detections lasting less than $\min(1/2,d/2)$ seconds are discarded. Here $d$ is the duration of the shortest event in the support set.
@@ 5.2 Comparison methods
We compare DRASDIC with essentially all of the previously published methods we are aware of for FSBSED that contain publicly available implementations. The first, “BEATS+linear” is a simple supervised baseline which consists of a frozen BEATS encoder [40] and a final linear layer. Support audio is windowed (4 seconds, 50% overlap) and the final layer is trained for 100 epochs to predict binary per-frame detection labels (final frame rate: 50 Hz). Training minimizes average per-frame binary cross-entropy loss. The initial learning rate of 0.01 (tuned using the validation set) is decayed to 0 using a cosine schedule. The second “AVES+linear” replaces the BEATS encoder with the pre-trained AVES encoder [41] (BirdAVES Base checkpoint), which was pre-trained on 2570 hours of animal sounds. “Protonet” is the prototypical network from [6], which itself adapts [7]. “Transductive” [9] uses a CNN encoder that is updated using unlabeled audio from the query set. “SCL” applies the supervised contrastive learning method introduced by [8]. “SCL+finetuning”, also introduced by [8] extends this by using support audio to fine-tune the encoder that was pre-trained using the SCL method.
@@
For Protonet, SCL, and SCL+Finetuning, we train a version using the training data from [4]. We also train a version using our generated scenes ($5e4$ scenes, each 40 seconds), which represents a $25\times$ increase in data quantity over the data used for training the original models.
@@
Figure 3: Qualitative results; events are in yellow. DRASDIC detects target sounds in dynamic environments (top two), but challenges include extremely low SNR (third), and the extended low-frequency drumming displays of ruffed grouse (bottom). Each spectrogram represents 10 seconds of audio.
5.3 Experiments
@@
We compare model performance on FASD13 (Table 2). Some datasets (JS, KD, MS, PB, PB24) contain events that are above DRASDIC’s 8kHz Nyquist frequency, or that are brief relative to the model’s 50Hz frame rate. For these, we give the model a 1/2-time version (1/6 for KD). We give other methods both the slowed and full-speed version of the data, and keep the version with the better score.
@@
On FASD13, DRASDIC outperforms all the alternatives on 6 out of 13 datasets. Across datasets DRASDIC has an average F1 score of .042 over the next best model. Compared to other methods that do not require gradient updates at inference time, DRASDIC outperforms the others on 9 of 13 datasets, and has an average F1 score of .173 over the next best model (64% relative improvement). Qualitatively, DRASDIC detected diverse target sounds, even amid others in the same frequency bands (Figure 3, top). Performance is strong across a variety of taxa and conditions. A failure case is for the JS dataset, which consists of jumping spider drumming. Here, the detection targets are specific drum types, and distinguishing between drum types relies partly on the rate of drumming. Our scene generator did not account for this type of information. Other failure cases are in Figure 3, bottom.
@@
For the comparison methods we trained with our generated data, there was no clear performance increase. These methods, which adopt a CNN architecture, employ a different few-shot mechanism than DRASDIC and also have fewer than $1/10$ the trainable parameters. The relative impact of these differences is investigated in Section 5.4.
5.4 Ablation experiments
In our main experiments, we scaled model parameters and data volume, while also adopting a few-shot mechanism that applied attention to support and query audio simultaneously. We conducted experiments to investigate the contributions of these changes, individually (Table 3, top). First, we compared our main model (116.1M parameters), whose transformer encoder has the same structure as BERT Base, to smaller versions based on BERT Small [42] (19.3M parameters) and BERT Tiny (2.5M parameters). Second, we compared our main data generation procedure to one that only generated 220 hours of unique scenes, and one that only generated 22 hours of unique scenes. Additionally, we compared to a version that used the non-synthetic training data (22 hours total) from [4], as well as a version for which 10% of the training examples were from [4] and the other 90% synthetic. Finally, we adjusted our few-shot mechanism to a prototypical network, which prevented attention from being applied to support and query audio simultaneously. For this, we kept the same architecture as our main method but applied a prototypical loss as in [6, 7]. Average performance on FASD13 and on the validation set dropped in all cases, indicating that each of these changes contributed to final model performance. Reducing data scale was especially damaging, likely due to the high number of trainable parameters in our main method.
@@ 5.4 Ablation experiments
We investigated the impacts of adjusting the randomness governing our scene generation procedure (Table 3, bottom). We perturbed the level of homogeneity of target events in a scene, the typical rate of events, the loudness of events, and whether we apply pitch shifting augmentations. Average performance is stable across some of these perturbations, but decreases when the randomness in event rate and event SNR is decreased. These parameters likely influence the level of diversity present across generated scenes more than the others.
@@
Eliminating random pitch shifts resulted in slightly better performance on FASD13. Designing a domain randomization strategy is an optimization problem, which we approached through a model selection criterion. This criterion did not produce the best model on the test set, aligning with the observation [6] that strong domain shifts between few-shot tasks present a challenge for FSBSED model development.
Table 3: Average F1 scores @0.3 IoU on FASD13 and validation [6] datasets. Model ablations appear on top, data ablations on bottom.
Method Avg (test) Avg (val)
DRASDIC .445 .704
BERT Small .425 .666
BERT Tiny .323 .521
Reduced data (220 h) .354 .504
Reduced data (22 h) .130 .088
Non-synthetic data (22 h) .135 .165
10% non-synthetic data .428 .666
Protonet loss .444 .638
High homogeneity in events .429 .630
Low homogeneity in events .444 .606
High events / second .381 .437
Low events / second .370 .669
Only high SNR events .402 .620
Only low SNR events .438 .682
No pitch/time shifting .457 .613